when the operation @ is defined, it means that P@Q=8P-Q. If A=3 and B=-2,what is A@B?
Answer by stanbon(75887)

when the operation @ is defined, it means that P@Q = 8P-Q. If A=3 and B=-2,what is A@B ?
--------------
3@-2 = 8*3-(-2) = 24+2 = 6
---------------
Cheers,
Stan H.
--------------